# FedLPPA: Learning Personalized Prompt and Aggregation for Federated Weakly-supervised Medical Image Segmentation

### v1

@@ III-D Preprocessing Operations for Bounding Box Labels
As stated in Sec. II-B, box annotations do not provide direct pixel-level supervision but rather serve as indicators of position and range. Hence, we transform them into alternative types of sparse labels for paradigm alignment. In Fig. 3, we present the preprocessing steps for the box annotations of the three modalities investigated in this paper: fundus image, OCTA image, and endoscopy image (the first two employ conventional boxes, while the latter utilizes rotated ones). After conversion, their ASP $p_{S,i}$ is set to match: medium sparsity for fundus images (converted to scribbles) and dense sparsity for the other two (converted to blocks). These preprocessing steps are specific to our experiments. Other methods, like using SAM-Med2D segmentation (boxes as prompts) [46] and moderate erosion to decrease label noise, are also viable.
@@ IV-A Datasets and Preprocessing
To verify the effectiveness of FedLPPA, we evaluate it on three segmentation tasks involving distinct modalities of medical imaging: segmenting optic disc/cup (OD/OC) from fundus images, segmenting foveal avascular zone (FAZ) from OCT angiography (OCTA) images, and segmenting polyps from endoscopy images. As stated in Table I, all fundus, OCTA and endoscopy images are acquired from various public datasets, possessing distinct data distributions even of the same modality. For fundus images, we follow Liu et al. [7] to crop the images around the OD area and resize them to 384$\times$384. For OCTA and endoscopy images, we directly resize them to 256$\times$256 and 384$\times$384. Data preprocessing includes normalizing all image intensities to [0,1], while data augmentation comprises random horizontal and vertical flips as well as random rotations (spanning from -45° to 45°).
@@
[…]
@@ D Kvasir-SEG[58] 900:100 Block
IV-B Implementation Details
[…]
@@ IV-C Comparisons with State-of-the-art
Tables III and IV respectively tabulate the quantitative results on the FAZ and Polyp segmentation tasks. As illustrated in Figs. 7B and 7C, unlike the fundus scenario, the distribution variations in OCTA and endoscopy images across different clients are relatively minor. Some pFL methods (like FedBN, FedAP, FedRep, and MetaFed), due to inappropriate or coarse-grained personalization, even underperform typical gFL methods like FedAvg and FedProx. Benefiting from the proposed TDF module which enables in-context learning for adaptive adjustments to different scenarios, and the PDLA mechanism, FedLPPA can make fine-grained personalized adjustments. Further, FedLPPA enables each client model to perceive different scenarios in an in-context learning manner and make finer-grained personalized adjustments. As a result, it achieves superior performance in various data heterogeneity tasks. Notably, in the FAZ segmentation task, FedLPPA significantly outperforms other FL methods and comes very close to the performance under the CT (full) setting. It is also worth mentioning that in the Polyp segmentation scenario, FedLPPA not only greatly surpasses its competitors but also exceeds the CT (full) setting by 4.6% in DSC, fully demonstrating the potential of WSS combined with FL. Visualization results for the three tasks are provided in Fig. 4.
@@ IV-D Ablation Studies and Analyses
TABLE V: Ablation study on the three key elements in FedLPPA. The columns for ODOC report the average metric values of OD and OC. TDF: tri-prompt dual-attention fusion, PD: prompt similarity dual-decoder, LA: learnable aggregation.
@@ IV-D1 Ablation Studies of FedLPPA
To delve deeper into FedLPPA, we perform ablations on the TDF module and the PDLA mechanism, where PDLA is divided into two parts: Prompt similarity Dual-decoder (PD, conducted on the server side) and Learnable Aggregation (LA, carried out on the client side). Table V presents the quantitative results of the main ablation studies across the three tasks. The first row represents the baseline, indicating a reduction to the classic FedAvg by removing the three components. The prompts in the TDF module are a prerequisite for acquiring the parameters of the auxiliary decoder. Hence, by initially integrating this module, a certain degree of performance improvement can be observed across all three tasks. In the primary ablation experiments, PD employs the Prompt similarity aggregated strategy mentioned in Sec. III-C. The inclusion of PD further significantly enhances performance on the previously established basis. The combination of the LA mechanism with TDF also yields commendable performance, but due to the absence of a dual-decoder to obtain reliable pseudo-proposals, it is not as effective as the TDF and PD combination on the ODOC and FAZ tasks (notably, a higher HD95 indicates a larger number of outlier predictions). With all three components integrated, the optimal performance is achieved, underscoring the indispensability of each component.
@@ IV-D2 Ablation Studies of TDF
In the ablation of the TDF module, we first retain the dual-attention fusion block and meticulously ablate the three types of prompts used in FedLPPA to explore the impact of each specific prompt. As DDP is used for measuring the data distribution of each client and forms the basis of the PD mechanism, it remains a constant element in this ablation experiment. From Table VI, it is observable that the addition of either ASP or UKP results in a certain improvement in performance, and the incorporation of both to form the triple prompt composition of FedLPPA achieves the optimal performance outcomes.
TABLE VI: Ablation study on the three prompts utilized in FedLPPA. DDP: data distribution prompt, ASP: annotation sparsity prompt, UKP: universal knowledge prompt.
@@
Given that the original triple prompts consist of only a limited number of bits, simply concatenating the prompts with sample features and processing them through several convolutional layers often proves inadequate for effective feature fusion and enhancement. Consequently, the effectiveness of the dual attention fusion block is further explored. Fig. 5 demonstrates that variants without any specific type of feature fusion nor enhancement achieve lower results in most metrics across the three tasks. Implementing either channel attention (CA) fusion or spatial attention (SA) fusion generally leads to performance improvements in most cases. However, exceptions are noted, such as that on the ODOC task where only using CA results in a performance decline. This could be attributed to the high data heterogeneity across different centers, where each local model is already capable of making personalized adjustments to various features, and the relatively coarse granularity of CA alone leads to a slight performance reduction. Yet, the application of dual attention fusion for feature enhancement consistently yields the best results in both types of metrics across the three tasks. This underscores the necessity of finer-grained modulation, fusion and enhancement of the triple prompts with sample features.
@@ IV-D3 Comparisons of Server-side Auxiliary Decoder Parameter Selection Strategies
In Sec. III-C, we present four strategies for obtaining the auxiliary decoder parameters on the server side, which will be distributed to the clients. These strategies include Random (RO), Fixed order (FO), Highest prompt similarity (HPS), and Prompt similarity aggregated (PSA, used as the default). In this subsection, we first explore the performance of these strategies and present the results in Fig. 6. The figure shows that the performance differences among the four strategies are not very significant across the three tasks. However, the PSA strategy consistently achieves the best performance in all scenarios. This is because the other three strategies only utilize information from a single separate client, while PSA aggregates and utilizes information from multiple similar sites based on prompt-calculated similarity. It also filters out information from dissimilar sites.
Figure 6: Performance of FedLPPA utilizing various server-side auxiliary decoder parameter selection strategies. RO, FO, HPS and PSA respectively denote random, fixed order, highest prompt similarity and prompt similarity aggregated strategies.
@@ IV-D4 Affinity Matrix Analysis
In Fig. 7, we perform t-SNE visualization (using the encoded features from the segmentation models trained in a centralized learning manner) for the three segmentation tasks, as well as the between-client similarity/affinity in each scenario (derived from our FedLPPA framework), calculated based on UKP and DDP. From the corresponding t-SNE and affinity matrices, strong correlation between data heterogeneity and affinity degree is observed. For example, in the ODOC task, where data heterogeneity among clients is high, the corresponding matrix shows lower similarity values and has zeros (indicating dissimilar clients). In contrast, in the FAZ and Polyp scenarios, where data heterogeneity is lower, the values in the affinity matrices are higher, especially in the Polyp task. The Polyp data has the least heterogeneity and all values in the corresponding affinity matrix are greater than 0.5. Furthermore, although not strictly corresponding, the relative distribution/relationship of data features in t-SNE and the similarities obtained from FedLPPA demonstrate a high degree of correlation, validating the effectiveness of using prompts to calculate similarities among multi-center data. Additionally, in conjunction with the results in Sec. IV-C, it clearly illustrates that FedLPPA effectively achieves personalization under various data heterogeneity scenarios, demonstrating its broad applicability, generalizability and robust performance.
Figure 7: The t-SNE visualization results of the correlation matrices of the three segmentation tasks. (A) ODOC, (B) FAZ and (C) Polyp.
@@ V Conclusion
This paper presents a novel personalized FL framework, FedLPPA, for tackling a practical but challenging federated heterogeneous WSS problem. A TDF module and a PDLA mechanism are introduced to enable in-context learning and fine-grained personalization. Extensive experiments are conducted on three medical image segmentation tasks, with superiority of FedLPPA being successfully established. Additionally, we have shown that under heterogeneous WSS conditions, FedLPPA can achieve segmentation performance that is comparable to that of labor-intensive fully-supervised centralized training.

### v2

@@ III-D Preprocessing Operations for Bounding Box Labels
As stated in Sec. II-B, box annotations do not provide direct pixel-level supervision but rather serve as indicators of position and range. Hence, we transform them into alternative types of sparse labels for paradigm alignment. In Fig. 4, we present the preprocessing steps for the box annotations of the four modalities investigated in this paper: fundus image, OCTA image, MRI and endoscopy image (the first three employ conventional boxes, while the latter utilizes rotated ones). After conversion, their ASP $p_{S,i}$ is set to match: medium sparsity for fundus images (converted to scribbles) and dense sparsity for the other three (converted to blocks). These preprocessing steps are specific to our experiments. Other strategies, like using SAM-Med2D for segmentation (boxes as prompts) [47] followed by moderate erosions to decrease label noise, are also viable.
@@ IV-A Datasets and Preprocessing
To verify the effectiveness of FedLPPA, we evaluate it on four segmentation tasks involving distinct modalities of medical imaging: segmenting optic disc/cup (OD/OC) from fundus images, segmenting foveal avascular zone (FAZ) from OCT angiography (OCTA) images, segmenting polyps from endoscopy images, and segmenting prostate from MRI. As stated in Table I, all fundus, OCTA, endoscopy and MRI are acquired from various public datasets, possessing distinct data distributions even of the same modality. For fundus images, we follow Liu et al. [7] to crop the images around the OD area and resize them to 384$\times$384. For OCTA and endoscopy images, we directly resize them to 256$\times$256 and 384$\times$384. For MRI, we utilize the multi-center dataset constructed by Liu et al. [48] for prostate MRI segmentation, dividing the cases into training and testing sets at a predetermined ratio and slicing them (slices not containing the prostate are excluded). The slices are then resized to 384$\times$384. Data preprocessing includes normalizing all image intensities to [0,1], while data augmentation comprises random horizontal and vertical flips as well as random rotations (spanning from -45° to 45°).
@@
IV-B Implementation Details
All compared FL methods and the proposed FedLPPA are implemented with PyTorch using NVIDIA A100 GPUs. We employ the vanilla UNet as the model architecture, with the number of channels progressively increasing from 16 to 256 from top to bottom. In FedLPPA, the auxiliary decoder shares the same architecture as the vanilla UNet decoder, and an MLP block is added to the main decoder. We optimize parameters using the AdamW optimizer with an initial learning rate of 1 $\times$ 10-2, and dynamically adjust the learning rate using a polynomial strategy with a power of 0.9. The trade-off parameter $\lambda$ utilized in $\mathcal{L_{\text{wss}}}$ (Eq. 15) and the batch size are respectively set to 0.5 and 12. We train all FL methods for 500 federation rounds to ensure a fair performance comparison. In each round of local updates, all clients undergo training for ten iterations, and in FedLPPA, two additional iterations are utilized to update $W_{M,i}$ and $W_{A,i}$ when $t>2$.
TABLE I: Details of the datasets utilized in our experiments. Scribble2 refers to scribbles of a different style generated using another algorithm.
@@ D Kvasir-SEG[60] 900:100 Block
MRI A NCI-ISBI 2013 (Siemens)[61] 340:81 Block 384$\times$384
B NCI-ISBI 2013 (Philips)[61] 297:87 Point
C I2CVB[62] 372:96 Scribble
D PROMISE12 (Siemens-UCL)[63] 134:41 Point
E PROMISE12 (GE-BIDMC)[63] 193:68 Scribble2
F PROMISE12 (Siemens-HK)[63] 122:36 Bounding box
@@ IV-C Comparisons with State-of-the-art
Tables III and IV respectively tabulate the quantitative results on the FAZ and Polyp segmentation tasks. As illustrated in Figs. 10B and 10C, unlike the fundus scenario, the distribution variations in OCTA and endoscopy images across different clients are relatively minor. Some pFL methods (like FedBN, FedAP, FedRep, and MetaFed), due to inappropriate or coarse-grained personalization, even underperform typical gFL methods like FedAvg and FedProx. Benefiting from the proposed TDF module which enables in-context learning for adaptive adjustments to different scenarios, and the PDLA mechanism, FedLPPA can make fine-grained personalized adjustments. Further, FedLPPA enables each client model to perceive different scenarios in an in-context learning manner and make finer-grained personalized adjustments. As a result, it achieves superior performance in various data heterogeneity tasks. Notably, in the FAZ segmentation task, FedLPPA significantly outperforms other FL methods and comes very close to the performance under the CT (full) setting. It is also worth mentioning that in the Polyp segmentation scenario, FedLPPA not only greatly surpasses its competitors but also exceeds the CT (full) setting by 4.6% in DSC, fully demonstrating the potential of WSS combined with FL. Table V tabulates the quantitative results for the prostate segmentation task. It is evident that FedLPPA continues to achieve the best overall performance, significantly surpassing CT (weak) and exhibiting the smallest disparity with CT (full), compared to other representative gFL and pFL methods. Representative visualization results are provided in Fig. 5 and Fig. 7.
@@
TABLE V: Performance comparisons of different FL methods as well as different local and centralized training settings on prostate segmentation. Note: The word ”Site” preceding the site letters is omitted to conserve table space.
DSC (Prostate) $\uparrow$ HD95 (Prostate) $\downarrow$
Methods A B C D E F Avg. A B C D E F Avg.
FedAvg[6] 83.83 81.93 78.16 80.79 66.47 82.12 78.88 10.33 28.49 31.29 10.77 36.23 17.01 22.35
FedProx[5] 83.03 82.76 78.47 82.70 58.70 84.54 78.37 11.56 28.88 26.42 10.08 34.52 15.34 21.13
FedBN[13] 86.58 79.83 78.16 77.93 74.31 87.01 80.64 8.65 94.83 17.64 74.84 26.24 10.09 38.72
FedLC[27] 84.07 80.96 77.09 82.39 73.10 79.17 79.46 10.40 26.20 24.85 10.42 31.28 24.79 21.32
FedALA[17] 85.52 84.84 81.21 83.78 71.32 87.95 82.44 9.36 24.98 18.22 15.06 32.02 9.41 18.18
FedICRA[16] 87.68 83.70 84.03 82.08 71.58 85.88 82.49 9.59 14.94 18.09 14.57 21.13 15.09 15.57
FedLPPA (Ours) 85.68 86.88 87.91 86.41 73.39 86.60 84.48 12.09 12.60 12.47 14.13 20.68 10.72 13.78
LT (Weak) 86.57 41.20 68.14 22.82 60.83 79.30 59.81 9.71 189.16 28.57 183.98 37.81 16.76 77.67
CT (Weak) 82.21 79.87 66.61 70.72 71.17 82.48 75.51 11.33 27.97 30.80 31.92 27.38 12.78 23.70
LT (Full) 90.92 88.36 89.93 73.96 86.10 91.48 86.79 7.99 11.35 9.42 40.52 20.79 7.92 16.33
CT (Full) 89.04 89.19 89.21 82.55 87.48 90.42 87.98 7.74 11.43 11.87 10.53 13.03 7.28 10.31
TABLE VI: Performance comparisons of typical FL methods using our dual-decoder WSS paradigm on FAZ and polyp segmentation. The values in parentheses indicate the differences in performance metrics compared to using the WSS paradigm with only sparse label for supervision; cyan denotes improvement, while orange denotes degradation.
Methods FAZ Polyp
[…]
FedAvg [6] 88.85 (+4.80) 10.44 (-3.74) 77.62 (+16.35) 67.22 (-8.43)
FedProx [5] 90.11 (+2.57) 8.17 (-0.17) 77.66 (+14.36) 60.45 (-5.50)
FedBN [13] 85.66 (+0.77) 9.51 (-1.76) 67.30 (+14.14) 51.87 (-53.44)
FedRep [12] 86.89 (+1.41) 12.51 (-1.09) 56.58 (+1.48) 80.91 (-3.67)
FedALA [17] 89.58 (+3.02) 8.64 (+0.74) 75.66 (+4.74) 67.50 (-4.98)
FedLPPA (Ours) 93.51 5.30 80.99 37.60
Considering that the communication efficiency is a crucial aspect of FL methods, we analyze this aspect by selecting several representative FL methods for comparative experiments with local iterations of 10, 20, 50, and 100 on the FAZ segmentation task—where a higher number of local iterations signifies lower communication frequency. The results, as displayed in Fig. 6, show that most FL methods exhibit performance degradation as communication frequency decreases. Notably, our FedLPPA exhibits the minimal performance decline and consistently outperforms all other methods at various communication frequencies.
To more fairly compare performance and underscore the effectiveness of our proposed framework, we implement the WSS paradigm used in FedLPPA for generating pseudo-proposal for several representative FL (including gFL and pFL) methods (employing the same set of loss functions, with the auxiliary decoder initialized differently from the main decoder as well as updated according to their respective aggregation methods or personalization strategies). We conduct experiments and performance comparisons on the FAZ and Polyp tasks, with results shown in Table VI. The results confirm that applying our WSS design to typical FL methods enhances performance to a certain extent, yet FedLPPA still maintains a substantial performance advantage.
Figure 6: Performance of representative FL methods under various local iterations for the FAZ segmentation task.
Figure 7: Visualization results from FedLPPA and other SOTA methods on prostate segmentation (CT denotes centralized training, with Weak and Full respectively indicating the use of sparse annotations and full masks for model training).
@@ IV-D Ablation Studies and Analyses
TABLE VII: Ablation study on the three key elements in FedLPPA. The columns for ODOC report the average metric values of OD and OC. TDF: Tri-prompt Dual-attention Fusion, PD: Prompt similarity Dual-decoder, LA: Learnable Aggregation.
@@ IV-D1 Ablation Studies of FedLPPA
To delve deeper into FedLPPA, we perform ablations on the TDF module and the PDLA mechanism, where PDLA is divided into two parts: Prompt similarity Dual-decoder (PD, conducted on the server side) and Learnable Aggregation (LA, carried out on the client side). Table VII presents the quantitative results of the main ablation studies across the ODOC, FAZ, and Polyp tasks. The first row represents the baseline, indicating a reduction to the classic FedAvg by removing the three components. The prompts in the TDF module are a prerequisite for acquiring the parameters of the auxiliary decoder. Hence, by initially integrating this module, a certain degree of performance improvement can be observed across all three tasks. In the primary ablation experiments, PD employs the Prompt Similarity Aggregated strategy mentioned in Sec. III-C. The inclusion of PD further significantly enhances performance on the previously established basis. The combination of the LA mechanism with TDF also yields commendable performance, but due to the absence of a dual-decoder to obtain reliable pseudo-proposals, it is not as effective as the TDF and PD combination on the ODOC and FAZ tasks (notably, a higher HD95 indicates a larger number of outlier predictions). With all components integrated, the optimal performance is achieved, underscoring the indispensability of each component.
@@ IV-D2 Ablation Studies of TDF
In the ablation of the TDF module, we first retain the dual-attention fusion block and meticulously ablate the three types of prompts used in FedLPPA to explore the impact of each specific prompt. As DDP is used for measuring the data distribution of each client and forms the basis of the PD mechanism, it remains a constant element in this ablation experiment. From Table VIII, it is observable that the addition of either ASP or UKP results in a certain improvement in performance, and the incorporation of both to form the triple prompt composition of FedLPPA achieves the optimal performance outcomes.
TABLE VIII: Ablation study on the three prompts utilized in FedLPPA. DDP: data distribution prompt, ASP: annotation sparsity prompt, UKP: universal knowledge prompt.
@@
Figure 9: Performance of FedLPPA utilizing various server-side auxiliary decoder parameter selection strategies. RD, FO, HPS and PSA respectively denote Random, Fixed Order, Highest Prompt Similarity and Prompt Similarity Aggregated strategies.
Given that the original triple prompts consist of only a limited number of bits, simply concatenating the prompts with sample features and processing them through several convolutional layers often proves inadequate for effective feature fusion and enhancement. Consequently, the effectiveness of the dual attention fusion block is further explored. Fig. 8 demonstrates that variants without any specific type of feature fusion nor enhancement achieve lower results in most metrics across the three (ODOC, FAZ and Polyp) tasks. Implementing either channel attention (CA) fusion or spatial attention (SA) fusion generally leads to performance improvements in most cases. However, exceptions are noted, such as that on the ODOC task where only using CA results in a performance decline. This could be attributed to the high data heterogeneity across different centers, where each local model is already capable of making personalized adjustments to various features, and the relatively coarse granularity of CA alone leads to a slight performance reduction. Yet, the application of dual attention fusion for feature enhancement consistently yields the best results in both types of metrics across the three tasks. This underscores the necessity of finer-grained modulation, fusion and enhancement of the triple prompts with sample features.
@@ IV-D3 Comparisons of Server-side Auxiliary Decoder Parameter Selection Strategies
In Sec. III-C, we present four strategies for obtaining the auxiliary decoder parameters on the server side, which will be distributed to the clients. These strategies include Random (RD), Fixed Order (FO), Highest Prompt Similarity (HPS), and Prompt Similarity Aggregated (PSA, used as the default). In this subsection, we first explore the performance of these strategies and present the results in Fig. 9. The figure shows that the performance differences among the four strategies are not very significant across the three tasks. However, the PSA strategy consistently achieves the best performance in all scenarios. This is because the other three strategies only utilize information from a single separate client, while PSA aggregates and utilizes information from multiple similar sites based on prompt-calculated similarity. It also filters out information from dissimilar sites.
PCA                 t-SNE          Correlation Matrix
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: The PCA, t-SNE visualization results, and the correlation matrices for three ((A) ODOC, (B) FAZ, and (C) Polyp) segmentation tasks.
@@ IV-D4 Affinity Matrix Analysis
In Fig. 10, we perform principal component analysis (PCA) and t-SNE visualization (using the encoded features from the segmentation models trained in a centralized learning manner) for three (ODOC, FAZ, and Polyp) segmentation tasks, as well as the inter-client similarity/affinity in each scenario (derived from our FedLPPA framework), calculated based on UKP and DDP. From the corresponding PCA, t-SNE and affinity matrices, strong correlation between data heterogeneity and affinity degree is observed. For example, in the ODOC task, where data heterogeneity among clients is high, the corresponding matrix shows lower similarity values and has zeros (indicating dissimilar clients). In contrast, in the FAZ and Polyp scenarios, where data heterogeneity is lower, the values in the affinity matrices are higher, especially in the Polyp task. The Polyp data has the least heterogeneity and all values in the corresponding affinity matrix are greater than 0.5. Furthermore, although not strictly corresponding, the relative distribution/relationship of data features in PCA, t-SNE and the similarities obtained from FedLPPA demonstrate a high degree of correlation, validating the effectiveness of using prompts to calculate similarities among multi-center data. Additionally, in conjunction with the results in Sec. IV-C, it clearly illustrates that FedLPPA effectively achieves personalization under various data heterogeneity scenarios, demonstrating its broad applicability, generalizability and robust performance.
IV-D5 Model and Communication Efficiency
In Table IX, we present the local model parameter size, communication overhead required per communication round, training time per federated round, and testing efficiency for various FL methods. It is important to note that these metrics for different FL methods are provided under identical hardware and communication conditions. Although FedLPPA shows a slight decrease in efficiency compared to other FL methods, its trade-off in efficiency is not significant relative to the performance improvement it offers.
TABLE IX: Comparison of model and communication efficiency. U: Upload, D: Download, N: Number of Clients.
Methods Model Parmas (M) Communication Overhead (M) Training Time (min/round) Test Efficiency (samples/s)
FedAvg 1.8135 [1.8135 (U) + 1.8135 (D)]×$\times N$ 0.200 43
FT 1.8135 [1.8135 (U) + 1.8135 (D)]×$\times N$ 0.200 43
FedProx 1.8135 [1.8135 (U) + 1.8135 (D)]×$\times N$ 0.724 43
FedBN 1.8135 [1.8105 (U) + 1.8105 (D)]×$\times N$ 0.204 43
FedAp 1.8135 [1.8105 (U) + 1.8105 (D)]×$\times N$ 0.654 43
FedRep 1.8135 [1.8132 (U) + 1.8132 (D)]×$\times N$ 0.202 43
MetaFed 1.8135 [1.8105 (U) + 1.8105 (D)]×$\times N$ 0.722 43
FedLC 2.1109 [2.1099 (U) + 2.1099 (D)]×$\times N$ 0.634 36
FedALA 1.8135 [1.8135 (U) + 1.8135 (D)]×$\times N$ 0.456 43
FedICRA 2.3372 [2.3372 (U) + 2.3372 (D)]×$\times N$ 0.582 30
FedLPPA 2.8978 [2.2564 (U) + 2.8978 (D)]×$\times N$ 0.710 29
@@ V Conclusion
This paper presents a novel pFL framework, FedLPPA, for tackling a practical but challenging federated heterogeneous WSS problem. A TDF module and a PDLA mechanism are introduced to enable in-context learning and fine-grained personalization. Extensive experiments are conducted on four medical image segmentation tasks, with superiority of FedLPPA being successfully established. Additionally, we have shown that under heterogeneous WSS conditions, FedLPPA can achieve segmentation performance that is comparable to that of labor-intensive fully-supervised centralized training.